OpenStudy (anonymous):

Arthur and Betty start walking toward each other when they are 100 m apart. Arthur has a speed of 3.0 m/s and Betty has a speed of 2.0 m/s. Their dog, Spot, starts by Arthur's side at the same time and runs back and forth between them at 5.0 m/s. By the time Arthur and Betty meet, what distance has Spot run?

OpenStudy (aaronq):

First find the time it took for Arthur and Betty to meet.  Use: \(x_f=x_i+v_0\Delta t+\dfrac{1}{2}a(\Delta t)^2\) Arthur: \(x_f=0+(3~m/s)\Delta t+\dfrac{1}{2}(0)(\Delta t)^2=(3~m/s)\Delta t\) Betty: \(x_f=100~m+(-2~m/s)\Delta t+\dfrac{1}{2}(0)(\Delta t)^2=100+(-2~m/s)\Delta t\) Their final position, \(x_f\),is the same. Make the equations equal to each other: \((3~m/s)\Delta t=100~m+(-2~m/s)\Delta t\) Solve for \(\Delta t\): \(\Delta t=\dfrac{100~m}{2~m/s+3m/s}=\dfrac{100~m}{5~m/s}=20~s\) Now we know the dog runs at 5 m/s, so if he ran for 20 s, he ran a total distance of: \(\Delta x=v_o\Delta t=(5~m/s)*20s=100 ~m\)